# Two prover perfect zero knowledge for MIP*

### v1

@@ 1. Introduction
Our results also have applications for the membership problem for quantum correlations. For exact membership, the cohalting problem is many-one reducible to membership in the set of quantum-approximable correlations $C_{qa}$, and to membership in the set of commuting operator correlations $C_{qc}$ [Slo19, CS19, FMS21]. It follows from $\operatorname{MIP}^{*}=\operatorname{RE}$ that the halting problem is Turing reducible to approximate membership in $C_{q}$, the set of quantum correlations, but this is not a many-one reduction. Theorem 1.1 immediately implies that there is a many-one reduction from the halting problem to approximate membership in $C_{q}$.
@@ 2. Nonlocal games and MIP*
There is a two-prover one round $\operatorname{AM}^{*}(2)$ protocol ($(\{\mathcal{G}_{x}\},S,V)$ for the halting problem with completeness $c=1$ and soundness $s=1/2$, such that $\mathcal{G}_{x}$ is a synchronous game with constant length questions, and $\operatorname{polylog}(|x|)$ length answers. Furthermore, if $\mathcal{G}_{x}$ has a perfect strategy, then it has a perfect oracularizable synchronous strategy.
@@ 4. BCS algebras and approximate representations
If $B=(X,\{(V_{i},C_{i})\}_{i=1}^{m})$ is a BCS, then we let $\mathcal{A}(B)$ denote the free product $\mathcal{A}(B):=\ast_{i\in[m]}\mathcal{A}(V_{i},C_{i})$. We let $\sigma_{i}:\mathcal{A}(V_{i},C_{i})\to\mathcal{A}(B)$ denote the natural inclusion of the $i$th factor, so $\mathcal{A}(B)$ is generated by the involutions $\sigma_{i}(x)$ for $i\in[m]$ and $x\in V_{i}$. Equivalently, $\mathcal{A}(B)$ is generated by the projections $\sigma_{i}(\Phi_{V_{i},\phi})$ for $i\in[m]$ and $\phi\in C_{i}$. To avoid clogging up formulas with symbols, we’ll often write $\Phi_{V_{i},\phi}$ instead of $\sigma_{i}(\Phi_{V_{i},\phi})$ when it’s clear what subalgebra $\mathcal{A}(V_{i},C_{i})$ the element belongs to. As with $\mathcal{A}(V,C)$, representations $\alpha$ of $\mathcal{A}(B)$ are in bijective correspondence with families of projective measurements $\{M^{i}_{\phi}\}_{\phi\in C_{i}}$, $i\in[m]$ via the relation $M^{i}_{\phi}=\alpha(\Phi_{V_{i},\phi})$. If ($(\{M^{i}_{\phi}\},\ket{v},\mathcal{H})$ is a synchronous commuting operator strategy for $\mathcal{G}(B,\pi)$, and $\alpha:\mathcal{A}(B)\to\mathcal{B}(\mathcal{H})$ is the representation with $\alpha(\Phi_{V_{i},\phi})=M^{i}_{\phi}$, then $a\mapsto\braket{v}{\alpha(a)}{v}$ is a tracial state on $\mathcal{A}(B)$. [^1] Here an (abstract) state on a *-algebra $\mathcal{A}$ is a linear functional $\tau:\mathcal{A}\to\mathbb{C}$ such that $\tau(1)=1$, $\tau(a^{*}a)\geq 0$ for all $a\in\mathcal{A}$, and $\tau(a^{*})=\overline{\tau(a)}$ for all $a\in\mathcal{A}$. A state is tracial if $\tau(ab)=\tau(ba)$ for all $a,b\in\mathcal{A}$, and faithful if $\tau(a^{*}a)>0$ for all $a\neq 0$. Conversely, if $\tau$ is a tracial state on $\mathcal{A}(B)$, then the GNS representation theorem implies that there is a synchronous commuting operator strategy $\mathcal{S}=(\{M^{i}_{\phi}\},\ket{v},\mathcal{H})$ such that $\tau(a)=\braket{v}{\alpha(a)}{v}$ where $\alpha$ is the representation corresponding to $\{M^{i}_{\phi}\}$. Note that the trace is faithful on the image of the GNS representation. As a result, synchronous commuting operator strategies for $\mathcal{G}(B,\pi)$ and tracial states on $\mathcal{A}(B)$ can be used interchangeably, and in particular $p\in C_{qc}$ if and only if there is a tracial state $\tau$ with $p(\phi,\psi|i,j)=\tau(\Phi_{V_{i},\phi}\Phi_{V_{j},\psi})$ for all $i$,$j$, $\phi$, and $\psi$. A tracial state is said to be finite-dimensional if its GNS representation has a finite-dimensional Hilbert space, so finite-dimensional tracial states on $\mathcal{A}(B)$ can be used interchangeably with synchronous quantum strategies for $\mathcal{G}(B,\pi)$, and $p\in C_{q}$ if and only if there is a finite-dimensional tracial state $\tau$ with $p(\phi,\psi|i,j)=\tau(\Phi_{V_{i},\phi}\Phi_{V_{j},\psi})$ for all $i$,$j$, $\phi$, and $\psi$. There is also a class of states, called the Connes-embeddable tracial states, with the property that $p\in C_{qa}$ if and only if there is a Connes-embbedable tracial state $\tau$ such that $p(\phi,\psi|i,j)=\tau(\Phi_{V_{i},\phi}\Phi_{V_{j},\psi})$ for all $i$,$j$, $\phi$, and $\psi$ [KPS18].
@@ 5. Homomorphisms between BCS algebras
In addition to looking at BCS games, we also want to consider transformations between constraint systems and the corresponding games. To keep track of how near-perfect strategies change, we introduce a notion of homomorphism for weighted algebras. Recall that if $\mathcal{A}$ is a *-algebra, then $a\geq b$ if $a-b$ is a sum of hermitian squares, i.e. there is $k\geq 0$ and $c_{1},\ldots,c_{k}\in\mathcal{A}$ such that $a-b=\sum_{i=1}^{k}c_{i}^{*}c_{i}$. Two elements $a,b\in\mathcal{A}$ are said to be cyclically equivalent if there is $k\geq 0$ and $f_{1},\ldots,f_{k},g_{1},\ldots,g_{k}\in\mathcal{A}$ such that $a-b=\sum_{i=1}^{k}[c_{i},d_{i}]$, where $[c,d]=cd-dc$. We say that $a\gtrsim b$ if $a-b$ is cyclically equivalent to a sum of squares. (For more background on these definitions, see see e.g. [KS08, Oza13]).
@@
If $B=(X,\{(V_{i},C_{i})\}_{i=1}^{m})$ is a BCS and $\pi$ is a probability distribution on $[m]\times[m]$, define $\mathcal{A}_{inter}(B,\pi)$ to be the weighted algebra $(\mathcal{A}(B),\mu_{inter})$, where $\mu_{inter}$ is defined from $\pi$ as in 5.3.
@@
Let $B=\left(X,\{(V_{i},C_{i})\}_{i=1}^{m}\right)$ be a BCS, and let $B^{\prime}=\left(X^{\prime},\{(W_{i},D_{i})\}_{i=1}^{m}\right)$ be a BCS with $X\subset X^{\prime}$, $V_{i}\subseteq W_{i}$ for all $1\leq i\leq m$, $W_{i}\cap W_{j}=V_{i}\cap V_{j}$ for all $1\leq i,j\leq m$, and for all $1\leq i\leq m$, $\phi\in V_{i}$ if and only if there exists $\psi\in W_{j}$ with $\psi|_{V_{i}}=\phi$. Then for any probability distribution $\pi$ on $[m]\times[m]$, the homomorphism
@@ 6. BCS algebras, subdivision and stability
Part of the point of the definition of subdivisions is that they preserve the synchronous algebra of the system.
@@
Let $B=\left(X,\{(V_{i},C_{i})\}_{i=1}^{m}\right)$ be a BCS, and let $B^{\prime}=\left(X,\{V_{ij},D_{ij}\}_{i,j}\right)$ be a subdivision. Let $\pi$ be a probability distribution on $[m]\times[m]$, and let $\pi_{sub}$ be the probability distribution defined from $\pi$ as above. Then $\operatorname{SynAlg}(B,\pi)\cong\operatorname{SynAlg}(B^{\prime},\pi_{sub})$.
@@
Because every pair of elements $x,y\in V_{i}$ belongs to some $V_{ij}$, we get an isomorphism
@@
Note that $\mu_{inter}$ is the same as the weight function of the algebra $\mathcal{A}_{inter}(B,\pi)$ defined in 5.4, except that it’s defined on $\mathcal{A}_{free}(B)$ rather than $\mathcal{A}(B)$. The weight function $\mu_{sat}$ comes from the defining relations for $\mathcal{A}(B)$, while $\mu_{clause}$ comes from the defining relations for $\mathcal{A}(B^{\prime})$, so $\mathcal{A}_{free}(B,B^{\prime},\pi)$ is a mix of relations from $\mathcal{A}_{inter}(B,\pi)$ and $\mathcal{A}_{inter}(B^{\prime},\pi)$. As mentioned previously, the context $V_{i}$ has an order inherited from $X$, and this is used for the order of the product when talking about $\Phi_{V_{i},\phi}$ and $\Phi_{V_{ij},\phi}$ in $\mathcal{A}_{free}(B)$. In particular, the order on $V_{ij}$ is compatible with the order on $V_{i}$.
The weight functions $\mu_{inter}$, $\mu_{sat}$ and $\mu_{clause}$ can also be defined on $\ast_{i=1}^{m}\mathbb{C}\mathbb{Z}_{2}^{V_{i}}$ using the same formula as in 6.6, and we use the same notation for both versions. The following lemma shows that we can relax $\mathcal{A}_{inter}(B,\pi)$ to $(\ast_{i=1}^{m}\mathbb{C}\mathbb{Z}_{2}^{V_{i}},\mu_{inter}+\mu_{clause})$, as long as $\pi$ is maximized on the diagonal.
@@
Let $B=\left(X,\{(V_{i},C_{i})\}_{i=1}^{m}\right)$ be a BCS, and let $\pi$ be a probability distribution on $[m]\times[m]$ that is maximized on the diagonal. Let $\mu_{inter}$ and $\mu_{sat}$ be the weight functions defined above with respect to $\pi$. Then there is an $O(t)$-homomorphism $\mathcal{A}_{inter}(B,\pi)\to(\ast_{i=1}^{m}\mathbb{C}\mathbb{Z}_{2}^{V_{i}},\mu_{inter}+\mu_{sat})$, where $t$ is the connectivity of $B$. Furthermore, if $B^{\prime}=\left(X,\{V_{ij},D_{ij}\}_{i,j}\right)$ is a subdivision of $B$, then there is an $M^{2}$-homomorphism $(\ast_{i=1}^{m}\mathbb{C}\mathbb{Z}_{2}^{V_{i}},\mu_{inter}+\mu_{sat})\to(\ast_{i=1}^{m}\mathbb{C}\mathbb{Z}_{2}^{V_{i}},\mu_{inter}+\mu_{clause})$, where $M=\max_{i}m_{i}$ is the maximum number of clauses $m_{i}$ in constraint $i$.
@@
$\alpha\left(\sum_{\begin{subarray}{c}1\leq i\neq j\leq m\\ x\in V_{i}\cap V_{j}\end{subarray}}\pi(i,j)h(\sigma_{i}(x)-\sigma_{j}(x))\right)$ $\leq\sum_{\begin{subarray}{c}1\leq i\neq j\leq m\\ x\in V_{i}\cap V_{j}\end{subarray}}\pi(i,j)\left(16\sum_{\mathclap{\varphi\in\mathbb{Z}_{2}^{V_{i}}\setminus C_{i}}}\Phi_{V_{i},\varphi}+16\sum_{\mathclap{\varphi\in\mathbb{Z}_{2}^{V_{j}}\setminus C_{j}}}\Phi_{V_{j},\varphi}+4h(\sigma_{i}(x)-\sigma_{j}(x))\right)$
@@
$\leq\sum_{i,j}\sum_{\phi\not\in D_{ij}}\frac{\pi(i,i)}{m_{i}^{2}}\left(\dfrac{\operatorname{poly}(K)}{\pi(i,i)}\operatorname{def}(\tau;\mu_{comm})+2\|\Phi_{V_{ij},\phi}\|_{\tau}\right)$
@@
Applying 6.9 and 6.8 yields the result. ∎

### v2

@@ 1. Introduction
Our results also have applications for the membership problem for quantum correlations. For exact membership, the cohalting problem is many-one reducible to membership in the set of quantum-approximable correlations $C_{qa}$, and to membership in the set of commuting operator correlations $C_{qc}$ [Slo19, CS19, FMS21]. It follows from $\operatorname{MIP}^{*}=\operatorname{RE}$ that the halting problem is Turing reducible to approximate membership in $C_{q}$, the set of quantum correlations, but this is not a many-one reduction. The proof of Theorem 1.1 immediately implies that there is a many-one reduction from the halting problem to approximate membership in $C_{q}$.
@@ Acknowledgements
2. Background on *-algebras
We recall some of the key concepts in the theory of *-algebras. See [Oza13, Sch20] for a more complete background. A complex *-algebra $\mathcal{A}$ is a unital algebra over $\mathbb{C}$ with an antilinear involution $a\mapsto a^{*}$, such that and $(ab)^{*}=b^{*}a^{*}$. We let $\mathbb{C}^{*}\langle X\rangle$ denote the free complex *-algebra generated by the set $X$. If $R\subseteq\mathbb{C}^{*}\langle X\rangle$, we let $\mathbb{C}^{*}\langle X:R\rangle$ denote the quotient of $\mathbb{C}^{*}\langle X\rangle$ by the two-sided ideal generated by $R$. If $X$ and $R$ are finite then we call $\mathbb{C}^{*}\langle X:R\rangle$ a finitely presented *-algebra.
A *-homomorphism $\phi:\mathcal{A}\to\mathcal{B}$ between *-algebras is an algebra homomorphism such that $\phi(x^{*})=\phi(x)^{*}$ for all $x\in A$. A *-representation of $\mathcal{A}$ is a *-homomorphism $\rho:\mathcal{A}\to\mathcal{B}(\mathcal{H})$ from $\mathcal{A}$ to the *-algebra of bounded operators on the Hilbert space $\mathcal{H}$. If $\mathcal{A}$ and $\mathcal{B}$ are *-algebras, and $\mathbb{C}^{*}\langle X:R\rangle$ is a presentation of $A$, then *-homomorphisms $\mathcal{A}\to\mathcal{B}$ correspond to homomorphisms $\phi:\mathcal{C}\langle X\rangle\to\mathcal{B}$ such that $\phi(r)=0$ for all $r\in R$. Thus, a *-representation is an assignment of operators to the elements of $X$ that satisfies the defining relations $R$.
If $\mathcal{A}$ is a *-algebra, then $a\geq b$ if $a-b$ is a sum of hermitian squares, i.e. there is $k\geq 0$ and $c_{1},\ldots,c_{k}\in\mathcal{A}$ such that $a-b=\sum_{i=1}^{k}c_{i}^{*}c_{i}$. A finitely presented *-algebra $\mathcal{A}$ is called archimedean if for all $a\in\mathcal{A}$ there exists a $\lambda>0$ such that $a^{*}a\leq\lambda 1$. The algebras we consider in this work are all archimedean. If $f:\mathcal{A}\to\mathbb{C}$ is a linear functional then $f$ is positive if $f(a)\geq 0$ whenever $a\geq 0$. A state on $\mathcal{A}$ is a positive linear functional $\tau:\mathcal{A}\to\mathbb{C}$ with $\tau(a^{*}a)\geq 0$ for all $a\in\mathcal{A}$, $\tau(1)=1$ and $\tau(a^{*})=\overline{\tau(a)}$ for all $a\in\mathcal{A}$. A state is tracial if $\tau(ab)=\tau(ba)$ for all $a,b\in\mathcal{A}$, and faithful if $\tau(a^{*}a)>0$ for all $a\neq 0$. A tracial state $\tau$ induces the trace norm $\|a\|_{\tau}:=\sqrt{\tau(a^{*}a)}$, also called the $\tau$-norm. Trace norms are unitarily invariant, meaning that $\|uav\|_{\tau}=\|a\|_{\tau}$ for all $a\in\mathcal{A}$, and all unitaries $u$ and $v$. An element $u\in\mathcal{A}$ is called unitary if $u^{*}u=1=uu^{*}$.
If $\rho:\mathcal{A}\to\mathcal{B}(\mathcal{H})$ is a *-algebra representation, then a vector $|v\rangle\in\mathcal{H}$ is cyclic for $\rho$ if the closure of $\rho(\mathcal{A})|v\rangle$ with respect to the Hilbert space norm is equal to $\mathcal{H}$. A cyclic representation of $\mathcal{A}$ is a tuple $(\rho,\mathcal{H},|v\rangle)$, where $\rho$ is a representation of $\mathcal{A}$ on $\mathcal{H}$ and $|v\rangle$ is a cyclic vector for $\rho$. If $\tau:\mathcal{A}\to\mathbb{C}$ is a positive linear functional on $\mathcal{A}$, then there is a cyclic representation $\rho_{\tau}$ of $\mathcal{A}$, called the GNS representation of $\tau$, such that $\tau(a)=\bra{\xi_{\tau}}\rho_{\tau}\ket{\xi_{\tau}}$ for all $a\in\mathcal{A}$. Two representations $\rho:\mathcal{A}\to\mathcal{B}(\mathcal{H})$ and $\pi:\mathcal{A}\to\mathcal{B}(\mathcal{K})$ of $\mathcal{A}$ are unitarily equivalent if there is a unitary operator $U:\mathcal{H}\to\mathcal{K}$ such that $U\rho(a)U^{*}=\pi(a)$ for all $a\in\mathcal{A}$. If $\tau$ is the state defined by $\tau(a)=\bra{\xi}\rho(a)\ket{\xi}$ for all $a\in\mathcal{A}$ and some cyclic representation $(\rho,\mathcal{H},\ket{\xi})$, then $(\rho,\mathcal{H},\ket{\xi})$ is unitarily equivalent to the GNS representation. A state $\tau$ is finite-dimensional if the Hilbert space $\mathcal{H}_{\tau}$ in the GNS representation $(\rho_{\tau},\mathcal{H}_{\tau},|\xi_{\tau}\rangle)$ is finite-dimensional. A state $\tau$ on $\mathcal{A}$ is called Connes-embeddable if there is a trace-preserving embedding of $\mathcal{A}$ into the ultrapower of the hyperfinite $\text{II}_{1}$ factor.
If $\mathcal{A}$ is a *-algebra then two elements $a,b\in\mathcal{A}$ are said to be cyclically equivalent if there is $k\geq 0$ and $f_{1},\ldots,f_{k},g_{1},\ldots,g_{k}\in\mathcal{A}$ such that $a-b=\sum_{i=1}^{k}[f_{i},g_{i}]$, where $[f,g]=fg-gf$. We say that $a\gtrsim b$ if $a-b$ is cyclically equivalent to a sum of squares. If $\tau$ is a tracial state on $\mathcal{A}$ then $\tau(c_{i}^{*}c_{i})\geq 0$ and $\tau([f_{j},g_{j}])=0$. Thus if $a\gtrsim b$ then $\tau(a)\geq\tau(b)$, and if $a$ and $b$ are cyclically equivalent then $\tau(a-b)=0$.
The *-algebras we use in this work are built out of the group algebras of the finitely presented groups
$\mathbb{Z}_{q}^{*V}=\langle V:x^{q}=1\rangle\text{ and }\mathbb{Z}_{q}^{V}=\langle V:x^{q}=1,xy=yx\text{ for all }x,y\in V\rangle.$
The group algebra $\mathbb{C}\mathbb{Z}^{*V}_{q}$ is the *-algebra generated by variables $x\in V$ with the defining relations from $\mathbb{Z}_{q}^{*V}$, along with the relations $x^{*}x=xx^{*}=1$ for all $x\in V$. Similarly $\mathbb{C}\mathbb{Z}_{q}^{V}$ is the *-algebra generated by variables $x\in V$ with the defining relations of $\mathbb{Z}_{q}^{V}$, along with the relations $x^{q}=x^{*}x=xx^{*}=1$ for all $x\in V$. Notice that $\mathbb{C}\mathbb{Z}_{q}^{V}$ is the quotient of $\mathbb{C}\mathbb{Z}_{q}^{*V}$ by the relations $xy=yx$ for all $x,y\in V$. If $\mathcal{A}$ and $\mathcal{B}$ are complex *-algebras, then we let $\mathcal{A}\ast\mathcal{B}$ denote their free product, and $\mathcal{A}\otimes\mathcal{B}$ denote their tensor product. Both are again complex *-algebras.
When working with $\mathbb{C}\mathbb{Z}_{q}^{V}$, a monomial in $V$ is an element of the form $\prod_{x\in V}x^{a_{x}}$, where $0\leq a_{x}<q$. We say that the monomial contains a variable $y\in V$ if $a_{y}>0$. The degree of a monomial is $\sum_{x}a_{x}$. If $\mathcal{A}_{1}$ and $\mathcal{A}_{2}$ are *-algebras for which we have a defined notion of monomial, then a monomial in $\mathcal{A}_{1}\otimes\mathcal{A}_{2}$ is an element of the form $v_{1}v_{2}$, where $v_{i}$ is a monomial in $\mathcal{A}_{i}$. The degree of $v_{1}v_{2}$ is the sum of the degrees of $v_{1}$ and $v_{2}$, and a variable $y$ is contained in $v_{1}v_{2}$ if $y$ is contained in $v_{1}$ or $v_{2}$. For instance, a monomial in $\mathbb{C}\mathbb{Z}_{q_{1}}^{V_{1}}\otimes\mathbb{C}\mathbb{Z}_{q_{2}}^{V_{2}}$ is an element of the form $\prod_{x\in V_{1}}x^{a_{x}}\cdot\prod_{y\in V_{2}}y^{b_{y}}$, where $0\leq a_{x}<q_{1}$ and $0\leq b_{y}<q_{2}$. Similarly, a monomial in $\mathcal{A}_{1}\ast\mathcal{A}_{2}$ is an element of the form $v_{1}\cdots v_{k}$, where $v_{j}$ is a monomial in $\mathcal{A}_{i_{j}}$ for all $1\leq j\leq k$, and $i_{j}\neq i_{j+1}$ for all $1\leq j<k$. In this case, the degree of $v_{1}\cdots v_{k}$ is the sum of the degrees of $v_{1},\ldots,v_{k}$, and a variable $y$ is contained in $v_{1}\cdots v_{k}$ if $y$ is contained in one of the monomials $v_{1},\ldots,v_{k}$. In any *-algebra where we have a defined notion of monomial, a polynomial is a linear combination of monomials.
A $C^{*}$-algebra $\mathcal{A}$ is a complex *-algebra with a submultiplicative Banach norm that satisfies the $C^{*}$ identity $\|aa^{*}\|=\|a\|^{2}$ for all $a\in\mathcal{A}$. Every $C^{*}$-algebra can be realized as a norm-closed *-subalgebra of the algebra of bounded operators $\mathcal{B}(\mathcal{H})$ on some Hilbert space $\mathcal{H}$. A $C^{*}$-algebra is a von Neumann algebra if it can be realized as a *-subalgebra of $\mathcal{B}(\mathcal{H})$ which is closed in the weak operator topology. More background on $C^{*}$-algebras and von Neumann algebras can be found in [Bla06].
@@ 3. Nonlocal games and MIP*
There is a two-prover one round $\operatorname{AM}^{*}(2)$ protocol ($(\{\mathcal{G}_{x}\},S,V)$ for the halting problem with completeness $c=1$ and soundness $s=1/2$, such that $\mathcal{G}_{x}$ is a synchronous game with constant length questions, and $\operatorname{polylog}(|x|)$ length answers. Furthermore, if $\mathcal{G}_{x}$ has a perfect strategy, then it has a perfect oracularizable synchronous quantum strategy.
@@ 5. BCS algebras and approximate representations
If $B=(X,\{(V_{i},C_{i})\}_{i=1}^{m})$ is a BCS, then we let $\mathcal{A}(B)$ denote the free product $\mathcal{A}(B):=\ast_{i\in[m]}\mathcal{A}(V_{i},C_{i})$. We let $\sigma_{i}:\mathcal{A}(V_{i},C_{i})\to\mathcal{A}(B)$ denote the natural inclusion of the $i$th factor, so $\mathcal{A}(B)$ is generated by the involutions $\sigma_{i}(x)$ for $i\in[m]$ and $x\in V_{i}$. Equivalently, $\mathcal{A}(B)$ is generated by the projections $\sigma_{i}(\Phi_{V_{i},\phi})$ for $i\in[m]$ and $\phi\in C_{i}$. To avoid clogging up formulas with symbols, we’ll often write $\Phi_{V_{i},\phi}$ instead of $\sigma_{i}(\Phi_{V_{i},\phi})$ when it’s clear what subalgebra $\mathcal{A}(V_{i},C_{i})$ the element belongs to. As with $\mathcal{A}(V,C)$, representations $\alpha$ of $\mathcal{A}(B)$ are in bijective correspondence with families of projective measurements $\{M^{i}_{\phi}\}_{\phi\in C_{i}}$, $i\in[m]$ via the relation $M^{i}_{\phi}=\alpha(\Phi_{V_{i},\phi})$. If ($(\{M^{i}_{\phi}\},\ket{v},\mathcal{H})$ is a synchronous commuting operator strategy for $\mathcal{G}(B,\pi)$, and $\alpha:\mathcal{A}(B)\to\mathcal{B}(\mathcal{H})$ is the representation with $\alpha(\Phi_{V_{i},\phi})=M^{i}_{\phi}$, then $a\mapsto\braket{v}{\alpha(a)}{v}$ is a tracial state on $\mathcal{A}(B)$. Conversely, if $\tau$ is a tracial state on $\mathcal{A}(B)$, then the GNS representation theorem implies that there is a synchronous commuting operator strategy $\mathcal{S}=(\{M^{i}_{\phi}\},\ket{v},\mathcal{H})$ such that $\tau(a)=\braket{v}{\alpha(a)}{v}$ where $\alpha$ is the representation corresponding to $\{M^{i}_{\phi}\}$. Note that the trace is faithful on the image of the GNS representation. As a result, synchronous commuting operator strategies for $\mathcal{G}(B,\pi)$ and tracial states on $\mathcal{A}(B)$ can be used interchangeably, and in particular $p\in C_{qc}$ if and only if there is a tracial state $\tau$ with $p(\phi,\psi|i,j)=\tau(\Phi_{V_{i},\phi}\Phi_{V_{j},\psi})$ for all $i$,$j$, $\phi$, and $\psi$. Finite-dimensional tracial states on $\mathcal{A}(B)$ can be used interchangeably with synchronous quantum strategies for $\mathcal{G}(B,\pi)$, and $p\in C_{q}$ if and only if there is a finite-dimensional tracial state $\tau$ with $p(\phi,\psi|i,j)=\tau(\Phi_{V_{i},\phi}\Phi_{V_{j},\psi})$ for all $i$,$j$, $\phi$, and $\psi$. Similarly, $p\in C_{qa}$ if and only if there is a Connes-embbedable tracial state $\tau$ such that $p(\phi,\psi|i,j)=\tau(\Phi_{V_{i},\phi}\Phi_{V_{j},\psi})$ for all $i$,$j$, $\phi$, and $\psi$ [KPS18].
@@ 6. Homomorphisms between BCS algebras
In addition to looking at BCS games, we also want to consider transformations between constraint systems and the corresponding games. To keep track of how near-perfect strategies change, we introduce a notion of homomorphism for weighted algebras.
@@
If $B=(X,\{(V_{i},C_{i})\}_{i=1}^{m})$ is a BCS and $\pi$ is a probability distribution on $[m]\times[m]$, define $\mathcal{A}_{inter}(B,\pi)$ to be the weighted algebra $(\mathcal{A}(B),\mu_{inter})$, where $\mu_{inter}$ is defined from $\pi$ as in 6.3.
@@
Let $B=\left(X,\{(V_{i},C_{i})\}_{i=1}^{m}\right)$ be a BCS, and let $B^{\prime}=\left(X^{\prime},\{(W_{i},D_{i})\}_{i=1}^{m}\right)$ be a BCS with $X\subset X^{\prime}$, $V_{i}\subseteq W_{i}$ for all $1\leq i\leq m$, and $W_{i}\cap W_{j}=V_{i}\cap V_{j}$ for all $1\leq i,j\leq m$. Suppose that for all $i\in[m]$, $\phi\in V_{i}$ if and only if there exists $\psi\in W_{i}$ with $\psi|_{V_{i}}=\phi$. Then for any probability distribution $\pi$ on $[m]\times[m]$, the homomorphism
@@ 7. BCS algebras, subdivision and stability
One of the first things we notice about subdivision is that strategies for $\mathcal{G}(B,\pi)$ can be lifted to strategies for the subdivided game.
@@
Let $B=\left(X,\{(V_{i},C_{i})\}_{i=1}^{m}\right)$ be a BCS, and let $B^{\prime}=\left(X,\{V_{ij},D_{ij}\}_{i,j}\right)$ be a subdivision. Let $\pi$ be a probability distribution on $[m]\times[m]$, and let $\pi_{sub}$ be the probability distribution defined from $\pi$ as above. The homomorphism $\alpha:\mathcal{A}(B^{\prime})\to\mathcal{A}(B)$ defined by $\sigma_{ij}(x)\mapsto\sigma_{i}(x)$ is a $1$-homomorphism $\mathcal{A}_{inter}(B^{\prime},\pi_{sub})\to\mathcal{A}_{inter}(B,\pi)$, and also induces an isomorphism $\operatorname{SynAlg}(B^{\prime},\pi_{sub})\cong\operatorname{SynAlg}(B,\pi)$.
Here $\sigma_{ij}(x)$ denotes the copy of $x$ in $\mathcal{A}(W_{ij},D_{ij})\subseteq\mathcal{A}(B^{\prime})$.
@@
Let $h(a)=a^{*}a$ denote the hermitian square of $a$ as in Lemma 7.2. By definition, $\alpha(\sigma_{ij}(x)-\sigma_{kl}(x))=\sigma_{i}(x)-\sigma_{k}(x)$. Hence
$\alpha\Bigl{(}\sum_{\begin{subarray}{c}ij\neq kl\\ x\in V_{ij}\cap V_{kl}\end{subarray}}\pi_{sub}(ij,kl)h(\sigma_{ij}(x)-\sigma_{kl}(x))\Bigr{)}$ $=\sum_{\begin{subarray}{c}ij\neq kl\\ x\in V_{ij}\cap V_{kl}\end{subarray}}\frac{\pi(i,k)}{m_{i}m_{k}}h(\sigma_{i}(x)-\sigma_{k}(x))$
$\leq\sum_{\begin{subarray}{c}i\neq k\\ x\in V_{i}\cap V_{k}\end{subarray}}\pi(i,k)h(\sigma_{i}(x)-\sigma_{k}(x)),$
since each variable $x\in V_{i}$ appears in at most $m_{i}$ subclauses $V_{ij}$. Hence $\alpha:\mathcal{A}_{inter}(B^{\prime},\pi_{sub})\to\mathcal{A}_{inter}(B,\pi)$ is a $1$-homomorphism.
To show that the synchronous algebras are isomorphic, observe that since every pair of elements $x,y\in V_{i}$ belongs to some $V_{ij}$, there is an isomorphism
@@
Note that $\mu_{inter}$ is the same as the weight function of the algebra $\mathcal{A}_{inter}(B,\pi)$ defined in 6.4, except that it’s defined on $\mathcal{A}_{free}(B)$ rather than $\mathcal{A}(B)$. The weight function $\mu_{sat}$ comes from the defining relations for $\mathcal{A}(B)$, while $\mu_{clause}$ comes from the defining relations for $\mathcal{A}(B^{\prime})$, so $\mathcal{A}_{free}(B,B^{\prime},\pi)$ is a mix of relations from $\mathcal{A}_{inter}(B,\pi)$ and $\mathcal{A}_{inter}(B^{\prime},\pi)$. As mentioned previously, the context $V_{i}$ has an order inherited from $X$, and this is used for the order of the product when talking about $\Phi_{V_{i},\phi}$ and $\Phi_{V_{ij},\phi}$ in $\mathcal{A}_{free}(B)$. In particular, the order on $V_{ij}$ is compatible with the order on $V_{i}$.
The weight functions $\mu_{inter}$, $\mu_{sat}$ and $\mu_{clause}$ can also be defined on $\ast_{i=1}^{m}\mathbb{C}\mathbb{Z}_{2}^{V_{i}}$ using the same formula as in 7.6, and we use the same notation for both versions. The following lemma shows that we can relax $\mathcal{A}_{inter}(B,\pi)$ to $(\ast_{i=1}^{m}\mathbb{C}\mathbb{Z}_{2}^{V_{i}},\mu_{inter}+\mu_{clause})$, as long as $\pi$ is maximized on the diagonal.
@@
Let $B=\left(X,\{(V_{i},C_{i})\}_{i=1}^{m}\right)$ be a BCS, and let $\pi$ be a probability distribution on $[m]\times[m]$ which is maximized on the diagonal. Suppose $\mu_{inter}$, $\mu_{sat}$ and $\mu_{clause}$ are the weight functions defined above with respect to $\pi$. Then there is an $O(t)$-homomorphism $\mathcal{A}_{inter}(B,\pi)\to(\ast_{i=1}^{m}\mathbb{C}\mathbb{Z}_{2}^{V_{i}},\mu_{inter}+\mu_{sat})$, where $t$ is the connectivity of $B$. Furthermore, if $B^{\prime}=\left(X,\{V_{ij},D_{ij}\}_{i,j}\right)$ is a subdivision of $B$, then there is an $M^{2}$-homomorphism $(\ast_{i=1}^{m}\mathbb{C}\mathbb{Z}_{2}^{V_{i}},\mu_{inter}+\mu_{sat})\to(\ast_{i=1}^{m}\mathbb{C}\mathbb{Z}_{2}^{V_{i}},\mu_{inter}+\mu_{clause})$, where $M=\max_{i}m_{i}$ is the maximum number of clauses $m_{i}$ in constraint $i$.
@@
$\alpha\Bigl{(}\sum_{\begin{subarray}{c}1\leq i\neq j\leq m\\ x\in V_{i}\cap V_{j}\end{subarray}}\pi(i,j)h(\sigma_{i}(x)-\sigma_{j}(x))\Bigr{)}$ $\leq\sum_{\begin{subarray}{c}1\leq i\neq j\leq m\\ x\in V_{i}\cap V_{j}\end{subarray}}\pi(i,j)\Bigl{(}16\sum_{\mathclap{\varphi\in\mathbb{Z}_{2}^{V_{i}}\setminus C_{i}}}\Phi_{V_{i},\varphi}+16\sum_{\mathclap{\varphi\in\mathbb{Z}_{2}^{V_{j}}\setminus C_{j}}}\Phi_{V_{j},\varphi}+4h(\sigma_{i}(x)-\sigma_{j}(x))\Bigr{)}$
@@
$\leq\sum_{i,j}\sum_{\phi\not\in D_{ij}}\frac{\pi(i,i)}{m_{i}^{2}}\left(\dfrac{\operatorname{poly}(K)}{\pi(i,i)}\operatorname{def}(\tau;\mu_{comm})+2\|\Phi_{V_{ij},\phi}\|^{2}_{\tau}\right)$
@@
Applying 7.9 and 7.8 yields the result. ∎